If (5,6) IS ON A LINE WITH THE SLOPE OF 4/3, GIVE THE COORDINATES FOR ANOTHER POINT ON THE LINE.
Answer by ptaylor(1894)

Let P(5,6)=P(x1,y1)
Now we know that the equation for this line is :
y-y1=m(x-x1) where m is the slope. So we have
y-6=4/3(x-5) or
3y-18=4x-20 and
4x-3y=2
if x=0, then y=-2/3 the coordinates for another point on the line
Hope this helps---Happy holidays----ptaylor